Find the domain and range Of the function
f(x)=x-1/x-5 ,x is not equal to 5
1
Expert's answer
2020-11-02T20:47:50-0500

"f(x)=\frac{x-1}{x-5};x\isin(-\infty,5)\cup(5,\infty)"

"f(x)=\frac{x-1}{x-5}=\frac{x-5+4}{x-5}=1+\frac{4}{x-5};\frac{4}{x-5}\not=0"

"f(x)\isin(-\infty,1)\cup(1,\infty)"

Answer domain-"(-\infty,5)\cup(5,\infty)" ;range -"(-\infty,1)\cup(1,\infty)"